Find the orthogonal canonical reduction of
the quadratic form Q = 3x2+ 2y2— 2.5 xy.
Also give its principal axes. Finally, draw a
rough sketch of the orthogonal canonical
reduction of Q = 4.
1
Expert's answer
2019-12-26T13:21:37-0500

Solution:


"\begin{pmatrix}
 3 & -1.25 \\
 -1.25 & 2
\end{pmatrix}"

For find the ortogonal canonical reduction we make the following determinant:



"\begin{vmatrix}
 3-\lambda & -1.25 \\
 -1.25 & 2-\lambda
\end{vmatrix} =0"

"\lambda^2-5\lambda+\frac{71}{16}=0"

This equation has irrational roots.


"\lambda_1=\frac{10+\sqrt{29}}{4}"


"\lambda_2=\frac{10-\sqrt{29}}{4}""f=\frac {10+\sqrt{29}}{4}y_1^2+\frac{10-\sqrt{29}}{4}y_2^2"

f is a ortogonal canonical reduction.

To go to the main axes, we solve the following systems of equations.


"\begin{alignedat}{2}
 (3-\lambda)&x_1- &1.25&x_2 = 0 \\
 -1.25&x_1+&(2-\lambda)&x_2 = 0
\end{alignedat}"

Next, "\lambda=\lambda_1" and "\lambda=\lambda_2" are considered.

This means that the given quadratic form is reduced to the principal axes by an orthogonal linear transformation.


"y_1=\frac{5}{\sqrt{58-4\sqrt{29}}}x_1+\frac{2-\sqrt{29}}{\sqrt{58-4\sqrt{29}}}x_2"

"y_2=\frac{5}{\sqrt{58+4\sqrt{29}}}x_1+\frac{2+\sqrt{29}}{\sqrt{58+4\sqrt{29}}}x_2"

For rough sketch of the orthogonal canonical reduction of Q = 4 need to sketch an ellipse given by equation:

"\frac{\tilde{x}^2}{\frac{16}{10-\sqrt{29}}}+\frac{\tilde{y}^2}{\frac{16}{10+\sqrt{29}}}=1"